OpenStudy (anonymous):

Factorize x^2+4x+4 I'm clueless. The formula, a^2+2ab+b^2=(a+b)^2, doesn't help either.

OpenStudy (cwrw238):

you need 2 numbers a and b where a* b = +4 and a + b = + 4 can you find them?

OpenStudy (cwrw238):

the factors will be in the form (x + a)(x + b)

OpenStudy (cwrw238):

a and b may be different values but also can be the same.

OpenStudy (anonymous):

So a and b = 2?

OpenStudy (cwrw238):

right

OpenStudy (cwrw238):

so the factors are (x + 2)(x + 2) or we can write that as (x + 2)^2

OpenStudy (anonymous):

Right, thanks. I think I probably get it now. Maybe. :P

OpenStudy (cwrw238):

lol - yw - practice some and it will stick. eg x^2 + 5x + 6 here ab = 6 and a + b = 5 what fits ?? a=2 and b = 3 so its (x + 2)(x + 3)

OpenStudy (anonymous):

So, a*b has to equal 6, and a+b has to equal 5. a=3 and b=2 fits. Therefore, the answer is (x+3)(x+2). Thanks. :) I'll come up with more similar problems and solve them. I think I get the fundamental train of thought and what not now. You've been most helpful, thank you very much! (:

OpenStudy (cwrw238):

yw